OpenStudy (kanwal32):

If (1+3+5+...+a)+(1+3+5+..+b)=(1+3+5.....c), where each set of parentheses contains the sum of consecutive odd integers as shown such that- (1) a+b+c=21 (2)a>6 if G=Max{a,b,c} and L=Min{a,b,c} then (A)G-L=4 (B)b-a=2 (C)G-L=7 (D)a-b=2

OpenStudy (kanwal32):

@ikram002p @hartnn @ganeshie8 hlp

OpenStudy (kanwal32):

@dan815

OpenStudy (kanwal32):

pls hlp

OpenStudy (kanwal32):

@hartnn hlp

OpenStudy (kanwal32):

y no 1 is replying?

ganeshie8 (ganeshie8):

cos it is a hard problem

OpenStudy (kanwal32):

yes it is

OpenStudy (kanwal32):

@satellite73

OpenStudy (kanwal32):

G-L=7 cannot be the answer cause odd-odd is even

ganeshie8 (ganeshie8):

have you gotten this far : \[\large (a+1)^2 + (b+1)^2 = (c+1)^2\] ?

OpenStudy (kanwal32):

ya

OpenStudy (kanwal32):

yes

ganeshie8 (ganeshie8):

the problem simplifies to solving pythagorean triples : \[\large x^2 + y^2 = z^2\] \(x+y+z = 24\) \(x \gt 7\)

OpenStudy (kanwal32):

ok

OpenStudy (kanwal32):

a,b,c will be max when a=b=c

OpenStudy (kanwal32):

G={7,7,7}

OpenStudy (kanwal32):

pls reply

OpenStudy (kanwal32):

@kainui

hartnn (hartnn):

x+y+z =24 x^2+y^2=z^2 only 6,8,10 satisfies this

hartnn (hartnn):

so, a,b,c = 5,7,9 which gives b-a=2

hartnn (hartnn):

Also, G= 9, L=5 so, G-L also = 4

OpenStudy (anonymous):

@hartnn great job man :) Genius

OpenStudy (anonymous):

=D

hartnn (hartnn):

ganeshie solved most of it!

OpenStudy (anonymous):

:)

OpenStudy (kanwal32):

@hartnn how to take G

OpenStudy (kanwal32):

and L

hartnn (hartnn):

G is max of (5,7,9) maximum among them is 9 so, G is 9

hartnn (hartnn):

L is min (5,7,9) minimum among them is 5, so L=5

OpenStudy (kanwal32):

OK

OpenStudy (anonymous):

awwww =D

OpenStudy (kanwal32):

Thnx @ganeshie8 and @hartnn

OpenStudy (anonymous):

:@

hartnn (hartnn):

lol, i just saw a>6 :P

hartnn (hartnn):

so, actually a=7, b=5, c= 9

hartnn (hartnn):

and a-b =2

hartnn (hartnn):

G-L still remains 4

OpenStudy (anonymous):

nice relationship :D

OpenStudy (kanwal32):

YES I WAS JUST THINKING THAT